### v1

# Accelerating Containerized Service Delivery at the Network Edge

Abstract
Efficient container image distribution is crucial for enabling machine learning inference at the network edge, where resource limitations and dynamic network conditions create significant challenges. In this paper, we present PeerSync, a decentralized P2P-based system designed to optimize image distribution in edge environments. PeerSync employs a popularity- and network-aware download engine that dynamically adapts to content popularity and real-time network conditions using a sliding window mechanism. PeerSync further integrates automated tracker election for rapid peer discovery and dynamic cache management for efficient storage utilization. We implement PeerSync with 8000+ lines of Rust code and test its performance extensively on both physical edge devices and Docker-based emulations. Experimental results show that PeerSync delivers a remarkable speed increase of 2.72$\times$, 1.79$\times$, and 1.28$\times$ compared to the Baseline, Dragonfly, and Kraken, respectively, while significantly reducing peak cross-network traffic by 90.72% under congested and varying network conditions.
Index Terms:
edge computing, container image distribution, P2P architecture, model inference.
I Introduction
Model inference at the network edge is becoming increasingly popular in applications such as edge-based video analytics, real-time inference in smart cities, and IoT-based predictive maintenance. To support these scenarios, containerization technologies, such as Docker, are widely used to enable the rapid deployment and management of workloads at scale [38, 44]. Container images, representing multi-layered filesystems, are not merely software artifacts but structured data that must be efficiently stored, distributed, and retrieved across diverse environments, ranging from cloud platforms to the network edge [13, 16, 34, 47, 2, 22].
Deploying a containerized application involves retrieving the necessary image layers from a container registry (e.g., Docker Hub), unpacking them, mounting them using a layered filesystem, and initiating the container entrypoint. These images, structured as a stack of layers representing filesystem segments, are stored and managed in centralized container registries. While these registries efficiently facilitate container image distribution in cloud environments, they encounter unique challenges in edge computing scenarios.
  • High latency. Distributed edge devices face substantial delays when fetching image layers over wide-area networks, leading to degraded responsiveness and increased cold-startup times for services [16, 8, 11].
  • Inefficiency under high concurrency. The non-parallel architecture of container registry uplinks often causes severe bottlenecks when handling concurrent image-pulling requests from multiple edge devices [11]. The growing size of AI models exacerbates this issue.
To address these limitations, peer-to-peer (P2P) architectures have been explored to accelerate container image distribution while alleviating the load on centralized registries [16, 44, 48, 19, 22, 26, 49, 23, 27, 2, 43]. P2P systems inherently offer distributed replication, parallel data transfers, and resilience against single points of failure. Although P2P-based solutions have shown promise in cloud environments, their effectiveness in edge settings is constrained by the unique characteristics of edge computing:
  • Limited network bandwidth. Unlike high-capacity data center networks, edge environments often operate with significantly constrained bandwidth, hindering efficient data exchange [13].
  • Low storage scalability. Portable edge devices, which typically rely on constrained storage media like SD cards or embedded flash memory, often have limited storage. Unlike cloud data centers equipped with technologies like Storage Area Networks [41], their scalability is inherently restricted. Over time, the accumulation of large AI-related images and other data, such as raw sensor outputs, can quickly deplete available storage space.
  • Weak stability under dynamic topology. The dynamic nature of edge networks introduces frequent connectivity changes, requiring robust mechanisms to adaptively manage edge devices and maintain consistency without relying on centralized components such as static trackers [20]. [^1] A tracker is used to identify which peers possess or request a given object in a P2P network. See [20] for more details.
While P2P-based solutions such as EdgePier [8], Kraken [43], and Starlight [11] address some of these issues and achieve great performance in most cases, they are still based on the traditional P2P architecture and consequently fall short in effectively balancing network efficiency, storage optimization, and adaptability to edge-specific conditions. More specifically, they lack fine-grained control mechanisms and rely on static trackers, making them unsuitable for fluctuating network conditions. Starlight [11] minimizes container startup latency by redesigning deployment protocols and storage formats but requires intrusive modifications to cloud registries, limiting its applicability. Moreover, existing solutions often neglect factors such as dynamic network quality, content popularity, and effective storage utilization, leaving room for improvement. For instance, our experiments (detailed in Sec. II) demonstrate that maintaining a single copy of any image layer within a local area network (LAN) can significantly enhance image-fetching speed without overloading local storage.
TABLE I: A comparison between our work and state-of-the-art implementations. ‘F.D.’ stands for ‘Fully Decentralized’, while ‘S.D.’ stands for ‘Semi Decentralized’.
Reference PeerSync Kraken [43] Dragonfly [2] Starlight [11] EdgePier [8] FID [23]
Architecture P2P (F.D.) P2P (S.D.) P2P (S.D.) C/S P2P (F.D.) P2P (S.D.)
P2P flavor Impl. from scratch Impl. from scratch Impl. from scratch N/A IPFS BitTorrent
Scenario Edge Cloud Cloud WAN Edge General
Tracker placement dynamic static static N/A static static
Open-sourced
To overcome these challenges, we propose PeerSync, a non-intrusive, P2P-based system tailored to the unique demands of edge computing. PeerSync constructs a fully decentralized P2P network spanning multiple edge devices across different LANs. It features a P2P image download engine that optimizes distribution by periodically calculating the required content pieces for each missing layer of a requested image, leveraging a scoring system that evaluates both content popularity and the quality of the network connection between peers and the requesting edge device. Unlike traditional P2P architectures that rely on manually configured trackers, PeerSync autonomously elects trackers based on real-time metrics, including bandwidth availability and resource utilization, ensuring resilience and performance. PeerSync also employs a selective deletion mechanism to optimize storage utilization, maintaining only essential image layers within each LAN. We compare PeerSync with state-of-the-art solutions in Table I. In summary, our main contributions are as follows:
  1. 1.
    We accelerate edge container image distribution with a tailored P2P architecture, featuring a high-performance download engine, fault-tolerant tracker with election ability for faster peer discovery, and collaborative dynamic space reclamation for efficient storage management.
  2. 2.
    We dynamically assign scores to peers based on runtime metrics like content popularity and network conditions using a sliding window mechanism, significantly enhancing the flexibility of image distribution in dynamic edge networks.
  3. 3.
    We implement PeerSync with 8000+ lines of Rust code, resulting in a statically linked binary of just 8.8 MB, making it easily deployable on edge devices.
  4. 4.
    We conducted extensive experiments to verify PeerSync’s performance under different network conditions. On average, PeerSync achieves 2.72x faster distribution than the Baseline, 1.79x faster than Dragonfly, and 1.28x faster than Kraken. Additionally, PeerSync reduces peak cross-network traffic by 90.72% under congested and variable network conditions.
II Motivation
In this section, we demonstrate the limitations of existing approaches using Kraken [43] and Starlight [11], motivating the need for a more efficient solution tailored to edge computing environments.
II-A Traditional P2P Approaches
P2P architectures, such as Napster [10], BitTorrent [1], and IPFS [9], provide decentralized mechanisms where nodes act as both consumers and providers. While these systems excel in distributed replication and data sharing, they suffer from a critical drawback: the lack of bandwidth awareness, which results in suboptimal use of network resources, particularly in edge environments.
Figure 1: An edge computing environment with 2 LANs and the corresponding observed results when using BitTorrent for image downloading.
To demonstrate this limitation, we conducted an experiment using Kraken [43]. As depicted on the left side of Fig. 1, our testbed consisted of two LANs connected via a 100 Mbps link. Each LAN included three Raspberry Pi (RPi) 4 Model B devices connected to a 1 Gbps switch. In LAN 1, two hosts seeded a large container image, while requests for the image originated from LAN 2. The experiment was repeated with 1, 2, and 3 local seeding peers in LAN 2. The results, shown in the right side of Fig. 1, highlight a significant inefficiency: despite having local peers available in LAN 2, approximately 10% of the data blocks were still fetched from remote peers in LAN 1, consuming over 95% of the transit network bandwidth. This behavior stems from BitTorrent’s peer selection strategy, which prioritizes global distribution over localized resource optimization. Consequently, the uplink bandwidth became saturated, negating the benefits of local peers and limiting the speed advantage of P2P systems in edge environments.
Tracker-based peer discovery is also vulnerable in dynamic edge environments given its centralized architecture. When a tracker-carrying node fails, it leads to fragmented networks, increased latency, and the requirement of manual intervention [12, 32, 25].
II-B Image Structure-based Approaches
Another class of approaches, exemplified by Starlight [11], attempts to optimize container startup time by analyzing the structures of the requested images. Starlight removes unnecessary components and decouples the download and execution processes to minimize startup latency. While effective for some use cases, these optimizations rely on runtime trace collection, which assumes that future executions will follow similar patterns and the two phases can be effectively parallelized. This creates a critical drawback when dealing with diverse and unpredictable workloads. Moreover, Starlight retains a reliance on centralized registries, limiting its effectiveness in dynamic, resource-constrained edge environments.
Figure 2: Timeline diagram of a large model inference task. Top: Lazy Loading (e.g., Starlight [11]): (1) Download components (PyTorch, cuDNN library, etc.), (2) Download the model, (3) Load the ML framework, (4) Perform inference. Bottom: P2P (e.g., PeerSync): (1) Download the full image, (2) Load the ML framework, (3) Perform inference.
As shown in Fig. 2, Starlight’s lazy loading strategy prioritizes downloading smaller components (e.g., libraries) before larger files like machine learning models. However, this approach provides only minimal speedup for inference tasks, as the critical delay stems from downloading large model files before computation can begin. For example, uncompressed CUDA-accelerated versions of Meta’s Llama 3.1 [15] (e.g., ai/meta-llama:3.1-8B-Instruct-cuda-12.6) require approximately 21.4 GiB of disk space. This image is dominated by two major components: models (15 GiB or 70.10%, consisting of only four files using Safetensors [3]) and frameworks (6.2 GiB or 28.97%, comprising PyTorch [35] and CUDA libraries). Trace-based methods like Starlight struggle with such storage footprints, as the bulk of the data resides in a few large files. In contrast, P2P-based approaches, such as the one proposed in this work, optimize how data is transmitted rather than focusing on when or what to download. By leveraging multiple peers, P2P systems can accelerate the distribution of large data blocks, making them more effective for workloads dominated by large model components.
II-C Insights
Both Kraken and Starlight illustrate shortcomings in existing approaches. Kraken, while leveraging P2P architectures, suffers from bandwidth inefficiencies and excessive reliance on centralized trackers. Starlight, despite its innovative lazy-loading approach, struggles with scalability across diverse model inference workloads and fails to address large model file distribution effectively. These limitations highlight the need for a decentralized, bandwidth-aware solution that maximizes local resource utilization, minimizes bandwidth waste, and adapts dynamically to the constraints of edge environments.
III System Design
III-A Overview
Fig. 3 illustrates the architecture of PeerSync, comprising four main components: Container API Interface (Sec. III-B), Download Engine (Sec. III-C), Embedded Tracker (Sec. III-D), and Cache Manager (Sec. III-E). Container API Interface provides a seamless and non-intrusive replacement for conventional Docker registries by implementing a suite of APIs. Download Engine coordinates the image retrieval process, combining traditional HTTP-based methods with innovative P2P-based approaches. Embedded Tracker maintains an autonomous and self-healing tracking system to ensure swarm integrity, eliminating reliance on centralized components. Cache Manager optimizes storage by managing cached images dynamically, prioritizing frequently used images and removing unused ones with minimal performance impact.
Figure 3: PeerSync architecture. The yellow boxes are PeerSync components.
Fig. 3 also depicts the lifecycle of an image within PeerSync, where red lines represent the retrieval process, and blue lines represent the removal process.
[…]
III-B Container API Interface
Container API Interface leverages the standardized APIs established by the Open Container Initiative (OCI). [^2] https://opencontainers.org/ Specifically, we adapt a subset of the Docker registry APIs to enable comprehensive container image retrieval and management while preserving simplicity for the user. This integration allows for seamless image handling, effectively masking the inherent complexity of the process. A critical aspect of this implementation is the management of image pulls by the Docker engine, which typically involves multiple HTTP requests: one to retrieve the image manifest and several more for each missing layer. The manifest, typically a JSON structure, contains essential details about the requested image, while layer requests return the actual data in the form of tarballs.
Our approach strategically optimizes these two API endpoints:
  • Manifest. Manifests are cached in an in-memory storage structure, which is continuously updated to reflect any upstream changes from Docker Hub.
  • Layer. This endpoint is a streamlined encapsulation that delegates the complexities of the multi-strategy image downloading process to Downloading Engine.
The API encapsulation layer within Docker Engine provides a simplified interface. Upon initialization, PeerSync queries the Docker Engine’s version via a UNIX domain socket and subsequently supplies runtime data to other system components. Essentially, Container API Interface acts as a bridge between Docker Engine and PeerSync’s internal mechanisms. Through this efficient encapsulation, PeerSync maintains precise control and monitoring of the Docker Engine, enabling a focus on optimizing internal operations.
III-C Downloading Engine
Downloading Engine serves as a core component in PeerSync, managing DHT network connections and facilitating popularity- and network-aware P2P downloading upon requests.
III-C1 Request Dispatcher
Request Dispatcher interfaces directly with Container API Interface (Sec. III-B). Upon receiving a download request, it evaluates the optimal strategy for image retrieval, deciding between downloads from a center registry and leveraging the P2P network. The process begins by assessing the size of the requested layer, using information stored in the manifest cache. Next, the Request Dispatcher determines whether the content is available within the nearby P2P network, taking into account time constraints during the search process.
P2P Downloading or Docker Hub? The efficiency of locating peers in the Distributed Hash Table (DHT) network is critical, as this process, which involves multiple network hops, can be time-consuming. Poor network conditions between peers, such as limited bandwidth or packet loss, can further extend the duration of peer discovery. In our context, such delays are unacceptable, as they negatively impact user experience by causing prolonged waiting times. To address this, we implement a time limit for peer aggregation. If image availability is confirmed within this timeframe, it indicates that image holders are present in nearby networks, making P2P downloading the preferred method. However, if the time limit expires without success, it suggests the absence of local image holders or their significant distance from the requesting node, prompting the system to switch to direct downloading from Docker Hub.
TABLE II: Size distribution of popular image layers. Each percentage represents the proportion of layers that are smaller than the specified threshold.
Thres. (B) % $<$ Thres. Thres. (B) % $<$ Thres.
128 1.64 1 Ki 29.27
8 Ki 41.45 512 Ki 47.78
4 Mi 57.38 32 Mi 76.81
256 Mi 97.19 605.73 Mi 100
Partial P2P Downloading. We analyzed the layer sizes of the top 100 popular images from Docker Hub as of July 2024, and as shown in Table II, the majority of these layers are relatively small, with a median size of just 1.03 MiB. Based on this observation, Request Dispatcher employs a partial P2P-based downloading strategy that balances efficiency and resource utilization by avoiding the lengthy processes typically associated with full DHT and tracker-based peer discovery [42]. Comprehensive discovery processes can take disproportionately longer than downloading directly from the central registry. However, relying exclusively on registry downloads for these smaller layers could lead to registry overload. To address this, our partial P2P method prioritizes multicast-based local peer discovery, bypassing the more time-consuming DHT and tracker-based methods. This adjustment significantly reduces both data transfer time and volume. If the requested layer is not found on the local network, Request Dispatcher defaults to direct downloads from Docker Hub. Any subsequent requests for the same layer within the local network are then fulfilled internally, ensuring rapid access while minimizing network congestion on the center registry. This approach effectively optimizes the trade-off between fast access and reduced registry load.
III-C2 Popularity- and Network-Aware Downloading
In PeerSync’s P2P downloading process, content is segmented into blocks, a standard practice that enables simultaneous downloading of different blocks from multiple peers, thereby maximizing bandwidth utilization [46]. When Request Dispatcher opts to download a particular image layer, it provides P2P Downloader with the layer details and relevant peer information, initiating the download process. Fig. 4 illustrates the P2P Downloader’s workflow, which consists of five key stages:
  1. 1.
    The process begins with the selection of a batch of blocks for the current download cycle.
  2. 2.
    Potential peers are sampled based on their dynamically updated scores, maintained by Peer Aggregator. This scoring system assesses peers’ suitability for the download task by considering both content popularity and network quality.
  3. 3.
    The peer with the highest score for each block is then selected to fulfill the download request, ensuring the optimal peer is chosen for each block.
  4. 4.
    A download request is sent to the selected peer.
  5. 5.
    Upon receiving a block, its integrity is verified. If the block fails validation, it is returned to the pending queue for re-download; otherwise, it is stored in the cache and delivered to the client.
Figure 4: Workflow of P2P downloading in PeerSync.
The cycle repeats with a new batch of blocks once the current one is successfully downloaded and validated.
TABLE III: Download times for an 8194.5 MiB image in a 10 Gbps LAN environment with different block sizes.
Block Size (MiB) #Blocks Download Time (s)
256 33 58
128 65 59
32 257 64
16 513 63
8 1025 87
Determining the Block Size. The methodology for determining block size and selecting peers for download requests is crucial to optimizing performance. The block size must be carefully balanced: blocks that are too large reduce parallelism by decreasing the number of blocks available for simultaneous download, while blocks that are too small increase protocol overhead and computational load due to a larger Merkle tree and more hash operations. Our experimental analysis, shown in Table III, used an 8194.5 MiB image in a 10 Gbps LAN environment. The results indicated that smaller block sizes significantly prolonged the download time. Based on these findings, we empirically determined the block size using the following guidelines, where $L_{b}$ denotes the block size and $L_{i}$ denotes the image size:
$L_{b}=\left\{\begin{array}[]{ll}L_{i}/256&\text{if }L_{i}\geq 1024\text{ MiB},\\ L_{i}/64&\text{if 256 MiB }\leq L_{i}<1024\text{ MiB},\\ L_{i}/16&\text{if 16 MiB}\leq L_{i}<256\text{ MiB},\\ L_{i}&\text{otherwise.}\\ \end{array}\right.$ (1)
For smaller images, which can be quickly downloaded from a single source, we consolidate the image into a single block to minimize communication overhead while still applying peer scoring algorithms.
Peer Selection Mechanism. Upon determining the optimal block size, PeerSync proceeds to the critical phase of peer selection. This phase significantly impacts system efficiency, influencing both the download time and the load on network links, particularly in congested conditions. PeerSync’s selection algorithm evaluates peers based on three criteria: network position, network speed, and content popularity. Each peer is assigned a score within the range of [0,100][0,100][ 0 , 100 ], reflecting its suitability for selection, with higher scores correlating to a greater probability of being chosen for content retrieval. In the following, we refer to the device initiating the content request as the client.
Network Position and Speed. The first criterion in PeerSync’s peer selection is network position—whether a peer resides in the same LAN as the client. Local peers are prioritized, receiving the maximum score of 100. For peers outside the LAN, PeerSync employs a speed-based scoring algorithm, where higher network speeds result in higher scores. This strategy favors local peers when available, and otherwise selects the fastest peers. Let $s_{p}^{t}$ denote the speed from peer $p$ to the client at time step $t$, and $\bar{s}^{t}$ the average speed from all peers to the client at time step $t$. Additionally, PeerSync maintains a sliding window $\mathcal{W}_{p}$, of size $L$, for recording historical transmission speeds from peer $p$ to the client. Similarly, $\mathcal{W}$ records the global average speeds. The speed $s_{p}^{t}$ is updated as:
$s_{p}^{t}=\frac{\sum_{s_{p}^{t^{\prime}}\in\mathcal{W}_{p}}s_{p}^{t^{\prime}}\cdot e^{L-t^{\prime}}}{\sum_{s_{p}^{t^{\prime}}\in\mathcal{W}_{p}}e^{L-t^{\prime}}},$ (2)
and the global average speed $\bar{s}^{t}$ is calculated as:
$\bar{s}^{t}=\frac{\sum_{\bar{s}^{t^{\prime}}\in\mathcal{W}}\bar{s}^{t^{\prime}}\cdot e^{L-t^{\prime}}}{\sum_{\bar{s}^{t^{\prime}}\in\mathcal{W}}e^{L-t^{\prime}}}.$ (3)
The network score $\text{net}_{p}^{t}$ of peer $p$ is then obtained as:
$\text{net}_{p}^{t}=s_{p}^{t}-\bar{s}^{t}.$ (4)
After re-scaling, $\text{net}_{p}^{t}$ is mapped to a value within [0,100][0,100][ 0 , 100 ].
Content Popularity. The popularity of an image layer affects its retrieval efficiency. Clients should avoid selecting peers storing rare images to optimize bandwidth usage. Let $\mathcal{P}^{t}$ denote the set of peers known to the client at time step $t$, and $\mathcal{I}_{p}^{t}$ the set of images possessed by peer $p$. We compute the popularity $\rho_{l}$ of layer $l$ as:
$\rho_{l}=\frac{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}\epsilon_{l}^{i}}{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}1},$ (5)
where $\epsilon_{l}^{i}=0$ if layer $l$ exists in image $i$, and 1 otherwise. The popularity score of peer $p$ is then given by:
$\text{pop}_{p}^{t}=100\times\left(1-\frac{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}e^{-\lambda\cdot\rho_{l}}}{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}1}\right).$ (6)
Custom Scoring and Softmax Distribution. Recognizing different user needs, PeerSync allows users to define custom scoring functions. The custom score of peer $p$ is denoted by $\text{cst}_{p}^{t}$. The final score (we also name it utility) of peer $p$, denoted by $U(p;t)$, is a weighted sum of network speed, content popularity, and the custom score:
$U(p;t)=\alpha\cdot\text{net}_{p}^{t}+\beta\cdot\text{pop}_{p}^{t}+\gamma\cdot\text{cst}_{p}^{t},$ (7)
where $\alpha$, $\beta$, and $\gamma$ are weights. PeerSync uses a softmax function over the scores to compute the probability distribution for peer selection:
$\Pr\{p;t\}=\frac{\exp(U(p;t))}{\sum_{p^{\prime}\in\mathcal{P}}\exp(U(p^{\prime};t))}.$ (8)
This ensures peers with higher scores are more likely to be selected.
Theoretical Analysis. Let $p_{t}$ be the peer selected by PeerSync and $p_{t}^{*}$ the peer that maximizes utility for a given client at time $t$. The instantaneous regret for selecting peer $p_{t}$ is defined as:
$R_{t}=U(p_{t}^{*};t)-U(p_{t};t).$ (9)
Theorem 1.
PeerSync’s peer selection mechanism achieves sublinear regret in terms of the cumulative difference between the optimal utility and the utility obtained by PeerSync. Specifically, over $T$ rounds of peer selection for different clients and content requests, the cumulative regret is bounded by $O(\sqrt{T})$, i.e.,
$R(T)=\sum_{t=1}^{T}(U(p_{t}^{*};t)-U(p_{t};t))\leq O(\sqrt{T}).$ (10)
The proof is omitted.
Proof.
The regret bound follows from the softmax-based selection strategy employed by PeerSync, where the probability of selecting a peer is a function of its utility score. The softmax function, with a decaying temperature parameter $\tau_{t}$, ensures exploration early on and focuses on exploitation as $\tau_{t}\to 0$, converging to the optimal peer.
The expected instantaneous regret $R_{t}$ is bounded by:
$\mathbb{E}[R_{t}]\leq\frac{C}{\tau_{t}},$ (11)
where $C$ is a constant representing the maximum possible utility difference between peers. Using $\tau_{t}=\frac{\tau_{0}}{\sqrt{t}}$, the cumulative regret over $T$ rounds is bounded by
$R(T)\leq\sum_{t=1}^{T}\frac{C}{\sqrt{t}}=O(\sqrt{T}),$ (12)
which means that PeerSync’s peer selection mechanism has a sublinear regret. ∎
III-D Embedded Autonomous Tracker
Trackers serve as centralized services crucial for facilitating peer discovery within P2P networks [21]. Without a functional tracker, peer discovery times can increase dramatically due to the multi-hop iteration of DHT, leading to inefficiencies. However, the centralized architecture of trackers poses potential points of failure, such as hardware malfunctions, software crashes, network outages, or inadequate maintenance. Additionally, managing a separate tracker infrastructure introduces administrative overhead.
To address these challenges, PeerSync incorporates an embedded autonomous tracker system that dynamically initiates a leader election process when no active trackers are available within the swarm. This system is triggered when a node detects the absence of live trackers or fails to connect to an existing tracker. The autonomous tracker supports the coexistence of multiple trackers but activates its election mechanism when all trackers are unavailable, particularly in local swarm regions. The node that identifies this situation initiates the election process using the FloodMax algorithm. [^3] FloodMax is a well-known leader election protocol in distributed systems. See [28] for more details. This algorithm enables each node to broadcast its stability metrics, such as uptime, to neighboring nodes. To mitigate the high communication overhead inherent in the FloodMax algorithm for large-scale networks, PeerSync implements path-pruning techniques [33]. These techniques reduce unnecessary message traffic, especially across network boundaries, achieving near-linear message complexity. This optimization minimizes resource consumption while ensuring the timely election of a new tracker.
Integrated directly into PeerSync’s codebase, Embedded Autonomous Tracker eliminates the need for manual configuration, significantly reducing the risk of single points of failure. By enhancing resilience and adaptability, this feature ensures uninterrupted operations in dynamic and distributed environments.
III-E Cache Manager
Efficient storage management is essential for maintaining the stability of low-end edge devices, which often operate with constrained resources [29]. Over time, PeerSync accumulates residual images in its cache, consuming valuable disk space, even when the associated applications are no longer active. Furthermore, upper-level applications require spare storage for data-heavy tasks, such as storing raw sensor metrics and datasets. Edge devices, typically reliant on flash-based storage without advanced virtualization technologies like Ceph [45], face additional limitations. Consequently, PeerSync must maintain its storage footprint as small as possible while preserving performance.
PeerSync addresses these challenges with a dynamic cache cleaner that leverages usage history to manage storage. This component monitors image usage through the Docker Engine, gathering statistics to determine which images to remove when storage reaches critical levels (e.g., less than 10% free space or a user-defined threshold). Image removal is prioritized based on three factors: last access time, image size, and content popularity (we consider both local and global popularity).
PeerSync’s storage management algorithm extends the classic Least Recently Used (LRU) cache replacement strategy by incorporating an additional metric: cache miss cost. Unlike traditional LRU, which assumes uniform memory access cost across an address space [39, 40], PeerSync accounts for the non-trivial time differences in retrieving an image:
  • Images present within the same LAN incur minimal retrieval cost, making them candidates for removal if redundancy exists.
  • Images that are the only copy within a LAN but have multiple copies elsewhere are prioritized for removal based on the number of external replicas.
  • Sole known copies of images are given the lowest removal priority, ensuring critical data remains accessible unless absolutely necessary.
This popularity-aware algorithm enables PeerSync to balance storage efficiency and performance, optimizing both retrieval times and disk utilization.
III-F Implementation Structure
PeerSync is designed with a message-passing architecture. Each component operates as an asynchronous task within the host system, with well-defined inter-component interfaces enabling communication and interaction. This approach provides both reliability and extensibility. In terms of reliability, isolated components ensure that failures do not affect the entire system. Any failed component can be quickly restarted to maintain continuous service. Extensibility is achieved through modular design, enabling easy updates to incorporate new features, adhere to evolving specifications, and support additional image formats.
PeerSync supports multiple peer discovery methods that function independently. This ensures Downloading Engine can immediately initiate fetching when an image is available from the P2P swarm, similar to BitTorrent. If P2P fetching is not feasible, PeerSync seamlessly falls back to direct HTTP downloading, ensuring uninterrupted operation.
TABLE IV: The container images chosen for evaluation.
Name Tag Service Compressed Size Description
redhat/granite-3-1b-a400m-instruct latest NLP 1.47 GB 1B finetuned IBM Granite 3.0 [18]
ai/meta-llama 3.1-8B-Instruct NLP 14.91 GB Llama 3.1 model by Meta [15]
cvisionai/segment-anything latest Vision 5.2 GB Segment Anything model by Meta [24]
langchain/langchain latest NLP 437.57 MB LLM application adaption framework[5]
pytorch/pytorch 2.5.1-cuda12.4-cudnn9-runtime General 3.11 GB Deep learning framework[35]
tensorflow/tensorflow nightly-gpu General 3.61 GB Deep learning framework[7]
IV Evaluation
We conduct extensive experimental evaluations in two environments: a Docker Compose-based emulation on a high-performance x86_64 host (Intel Xeon Silver 4214) and physical edge devices (Raspberry Pi 4 Model B). The emulation enables the simulation of large-scale environments, while the physical setup validates the practical feasibility of PeerSync in real-world scenarios. We select a range of popular AL/ML and commonly used container images of varying sizes, from small base images to large language model (LLM) containers, as listed in Table IV. These images represent diverse use cases, allowing us to evaluate system performance under varying workloads and conditions.
We evaluated PeerSync against two popular container image distribution systems: Dragonfly [2] and Kraken [43], as well as the conventional HTTP-based pull method, referred to as the Baseline. Dragonfly, hosted by the Cloud Native Computing Foundation (CNCF) as an Incubating Level Project, [^4] https://d7y.io/ leverages peer-to-peer (P2P) technology to accelerate file distribution and container image delivery. It is designed to enhance efficiency in distributed environments by reducing reliance on centralized registries. Kraken, an open-source project developed by Uber and deployed in production since early 2018, [^5] https://github.com/uber/kraken also employs P2P technology for Docker image management, replication, and distribution. Kraken is particularly tailored for hybrid cloud environments, offering decentralized capabilities to optimize large-scale image deployments. Although Starlight [11] also focuses on improving container image delivery, it relies on fundamentally different mechanisms, such as runtime trace collection and dependence on centralized registries. These differences make direct comparisons with PeerSync less relevant (see Sec. II). Consequently, we did not include Starlight in our comparisons.
Figure 5: Average AI/ML image distribution time under various network conditions (log scale). The $x$-axis is the parameter $A$. The missing points mean that the download operation could not be completed during the limited time frame.
IV-A Docker Compose-based Emulation
IV-A1 Experimental Setup
We create 10 Linux bridge networks ($n$}, where $n=1$ to 101010) to emulate a distributed environment with interconnected routers. Each network contained a router for inter-network communication and seven worker nodes requesting container images. The routers were configured with bandwidth limits (50 Mbps to 1 Gbps), latency, and packet loss using tc(8) [6], simulating real-world ISP and transit networks. Intra-network bandwidth was set to 1 Gbps with zero packet loss.
To model real-world conditions, we vary model inference request frequencies based on the following equation:
$t_{i}\sim\text{Poisson}^{-1}\left(\text{random}\left(0.001,A\cdot e^{\frac{B}{\text{s}_{i}}}\right)\right),$
where $\text{s}_{i}$ is the image size, and $A$ and $B$ are tunable parameters. Higher values of $A$ and $B$ increased the frequency of image-pulling requests, simulating high-demand scenarios. Background traffic generated by iPerf [4] created additional congestion, reflecting the shared link usage typical in real-world applications. All centralized components, including registries, Dragonfly databases, and Kraken trackers, were placed in net_worker1 to simulate centralized cloud environments.
IV-A2 Image Distribution Time
TABLE V: Average Speedup under high request frequency compared to the Baseline (smaller is better).
Net. Profile Baseline Dragonfly[2] Kraken[43] PeerSync
Stable 100% 75.01% 80.13% 77.85%
Congested 100% 64.27% 61.89% 45.89%
Varying 100% 65.63% 46.96% 36.71%
Fig. 5 illustrates the average container image distribution times under different network conditions, providing critical insights into the importance of local data prioritization. Since the Baseline method shows significant performance degradation as request frequency increases, a logarithmic scale (base 222) is used on the $y$-axis for clarity.
Stable network conditions (see top of Fig. 5). In environments with sufficient and stable network resources, all P2P-based methods, including PeerSync, exhibit similar performance trends as request frequency increases (see Table V). While distribution latency increases slightly with higher request loads, P2P methods consistently outperform the Baseline HTTP approach. This aligns with the known benefits of P2P models, such as efficient data sharing and reduced reliance on centralized registries, which have been extensively validated in data center contexts. The trend is particularly pronounced for large LLM images: the average distribution time roughly quadruples as request frequency increases, highlighting the critical role of P2P technologies in facilitating the deployment of large modern AI/ML models with billions of parameters.
Congested network conditions (see lower-left of Fig. 5). In environments with bandwidth limitations, packet loss, and high latency, the Baseline method suffers severe performance degradation due to the long-tail effects of centralized pulling. Remote data transfers are particularly affected, often exceeding the default time limit of 1200 seconds, as indicated by the missing points in the green Baseline lines. In contrast, PeerSync and other P2P-based methods maintain better performance by distributing the load across multiple peers, effectively mitigating bandwidth constraints. Notably, Dragonfly shows a sharp performance drop for larger images due to its reliance on centralized components such as schedulers and managers. As communication between edge nodes and these components becomes unreliable, Dragonfly’s ability to coordinate cloud-edge interactions weakens, resulting in bottlenecks during image serving. PeerSync, by contrast, leverages its decentralized architecture to achieve consistent performance, even under congestion.
Variable network conditions (see lower-right of Fig. 5). This scenario simulates real-world edge environments where latency, packet loss, and bandwidth fluctuate significantly, and participating nodes frequently join and leave the network. Such instability disrupts inter-peer connections and hinders image fetching, even for smaller images like langchain/langchain. Compared to other P2P-based solutions, PeerSync demonstrates superior resilience under these conditions. By prioritizing local peer synchronization and minimizing dependence on centralized components, PeerSync reduces reliance on specific, potentially unstable links and dynamically explores all available links to maximize performance. This strategy not only improves distribution efficiency but also enhances fault tolerance, which is critical in environments with unpredictable network behavior.
PeerSync’s ability to optimize local storage and peer communication makes it particularly robust in dynamic edge environments. As shown in Table V, PeerSync outperforms the Baseline, Dragonfly, and Kraken under unstable, high-frequency scenarios by average factors of 2.72, 1.79, and 1.28, respectively. [^6] In our experiments, the time limit for each download request was set to 1200 seconds. In some instances, the Baseline was unable to finish within this time frame, resulting in missing data points when calculating the average speedup. We assigned a value of 1200 seconds to the missing data points. Consequently, the actual speedup relative to the Baseline is higher than the calculated value. The speedup is even more pronounced for larger container images, reinforcing PeerSync’s effectiveness in addressing the challenges of large-scale image distribution.
IV-A3 Uplink Occupancy
Tables VI, VII, and VIII illustrate the cross-network traffic for each image distribution method across three distinct network profiles: stable, congested, and unstable conditions. Cross-network traffic represents data that traverses routers and encounters bandwidth limits, latency, and packet loss, effectively simulating the challenges of data traversing upstream ISP transit networks and the Internet. Unlike local network bandwidth, which is often abundant, cross-network bandwidth is typically constrained by external factors such as physical link limitations and tariff plans. Thus, minimizing cross-network traffic is critical, particularly in edge computing, where the goal is to prioritize intra-LAN resources as much as possible.
TABLE VI: Sum of cross-network traffic under sufficient and stable network conditions.
Solution Maximum (Gbps) Average (Gbps)
PeerSync 5.07 0.82
Kraken [43] 6.61 0.93
Dragonfly [2] 8.80 1.37
Baseline 11.91 6.18
Sufficient and stable network conditions. As shown in Table VI, PeerSync consistently minimizes cross-network bandwidth usage under stable conditions. By prioritizing local peer data, PeerSync achieves an average bandwidth consumption of only 0.82 Gbps, significantly lower than Kraken (0.93 Gbps), Dragonfly (1.37 Gbps), and the Baseline (6.18 Gbps). While Dragonfly and Kraken leverage P2P techniques, their reliance on centralized components for coordination and scheduling results in higher cross-network traffic compared to PeerSync. The Baseline method incurs the highest bandwidth usage due to its centralized pulling mechanism, which does not take advantage of local data sharing.
TABLE VII: Sum of cross-network traffic under congested but stable network conditions.
Solution Maximum (Gbps) Average (Gbps)
PeerSync 11.50 0.91
Kraken [43] 13.76 1.18
Dragonfly [2] 11.94 4.69
Baseline 10.36 9.81
Congested but stable network conditions. In congested but stable network conditions (Table VII), PeerSync continues to outperform other methods by maintaining an average cross-network traffic of just 0.91 Gbps. The Baseline approach shows the lowest maximum bandwidth (10.36 Gbps), which is a result of its slower and less efficient data retrieval. However, it saturates network capacity with an average bandwidth usage of 9.81 Gbps, underscoring its inefficiency in managing congested networks. Kraken and Dragonfly perform better than the Baseline but still exhibit performance degradation due to their centralized dependencies. Dragonfly, in particular, suffers from its reliance on schedulers and managers, resulting in an average traffic of 4.69 Gbps, over five times higher than PeerSync.
TABLE VIII: Sum of cross-network traffic under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions.
Solution Maximum (Gbps) Average (Gbps)
PeerSync 9.15 0.76
Kraken [43] 10.99 0.89
Dragonfly [2] 10.13 4.01
Baseline 8.87 6.34
Congested and unstable network conditions. Table VIII highlights the performance of each method under congested and unstable network conditions, characterized by fluctuating packet loss, latency, and bandwidth. PeerSync demonstrates remarkable robustness, maintaining the lowest average cross-network traffic at 0.76 Gbps. Kraken performs slightly worse, with an average traffic of 0.89 Gbps, but still manages to leverage its decentralized architecture to reduce reliance on external networks. Dragonfly, on the other hand, suffers significant performance degradation due to its dependence on centralized coordination, resulting in an average traffic of 4.01 Gbps. The Baseline method experiences the lowest maximum bandwidth usage (8.87 Gbps) but continues to exhibit high average traffic (6.34 Gbps), reflecting its inability to adapt to varying network conditions. This highlights the inefficiency of centralized pulling in handling unstable and dynamic environments.
The results across all network profiles clearly demonstrate PeerSync’s superior ability to minimize cross-network bandwidth consumption. This efficiency is achieved by leveraging local resources and reducing reliance on external network traffic, which is particularly critical in edge environments where bandwidth is constrained and often unstable.
TABLE IX: Number of edge devices in a LAN and the average image distribution time (over 100 image retrieval requests) provided by LAN cache.
Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
Average image distribution time (s) 9.11 8.33 9.69 11.34 8.87 8.06 5.08 2.86 2.89 2.19
TABLE X: Number of nodes (1 $\sim$ 5) and comparison of the sum of cache space occupied by Cache Cleaner and the LRU policy.
Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
Space occupied by Cache Cleaner (MiB) 196 388 475 589 593 668 905 1259 1199 1420
Space occupied by LRU (MiB) 110 206 411 701 837 948 973 1164 1056 1701
Figure 6: Average image distribution time under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions (log scale). The $x$-axis is the parameter $A$. (Smaller images)
IV-A4 Cache Strategy
We evaluated the Cache Cleaner algorithm by analyzing its effectiveness in managing requests and accelerating image retrieval within a LAN. Table IX shows the relationship between the number of edge devices in a LAN and the average time to download images. Initially, as the number of nodes increases, image distribution times rise slightly due to simultaneous requests to the upstream registry. However, as the LAN becomes more populated, the local cache network is increasingly leveraged, resulting in significantly reduced average retrieval times.
As nodes in the LAN form a collaborative cache, the Cache Cleaner algorithm effectively retains images that are likely to benefit the entire network. This approach optimizes local data sharing and minimizes redundant requests to external networks. The results demonstrate that as the cache network scales, the average download time decreases dramatically, underscoring the efficiency of PeerSync’s collaborative caching mechanism.
Comparison with the LRU policy. To further validate the Cache Cleaner’s performance, we compared its total cache usage against the Least Recently Used (LRU) policy in Table X. Cache Cleaner consistently achieves better space utilization by coordinating cache management across neighboring peers. This approach avoids redundant caching of identical content across multiple nodes, enabling more efficient storage utilization within the LAN. In contrast, LRU operates independently on each node and lacks collaboration, leading to higher total cache consumption. Cache Cleaner’s ability to prioritize fast intra-LAN retrieval not only optimizes storage but also improves overall system efficiency.
Distribution of smaller images. Although our primary focus is on large AI/ML images, smaller, non-AI base images are also widely used in practical setups. To emulate the operation of Docker Hub, which serves popular images in large volumes daily, we selected the 10 most downloaded images from Docker Hub and increased the request frequency by adjusting the parameter $A$. Fig. 6 presents the results, demonstrating PeerSync’s ability to maintain its performance advantage over other P2P-based solutions, even under high-frequency request conditions for smaller images. These results highlight PeerSync’s capability to handle the significant volumes typical of Docker Hub operations, proving its versatility across diverse workloads.
The experiment results validate the efficiency of PeerSync’s Cache Cleaner algorithm in optimizing local caching, reducing redundant external requests, and improving retrieval times within a LAN. Its collaborative caching strategy outperforms traditional policies like LRU, offering better space utilization and accelerated data distribution. Additionally, PeerSync demonstrates its ability to efficiently handle high-frequency requests for smaller images, further reinforcing its robustness and scalability across varied workloads and network conditions. This makes PeerSync an efficient solution for both edge and cloud-based environments.
IV-B Real-World Experiments
We deployed a physical testbed using Raspberry Pi devices to compare PeerSync’s performance against the Baseline, Dragonfly, and Kraken. As shown in Fig. 7, the setup comprises six Raspberry Pi (RPi) 4 Model B devices connected via two layer 2 switches, each with a 1 Gbps link speed. Each LAN contains three RPis, with the two switches connected to a common router. The router is configured to forward packets between the networks while limiting the inter-LAN bandwidth to 100 Mbps. To simplify deployment, static routes are configured on the router. This setup is similar to the Docker Compose-based emulation described earlier, with two key differences: a) Physical hardware and real-world networking scenarios are used; b) The aarch64 version of PeerSync is deployed instead of the amd64 version to accommodate the ARM-based Raspberry Pi architecture.
Figure 7: The testbed for real-world experiments.
Benchmark requests were made to agents running on each RPi to measure image distribution times. Table XI presents the 90th and 100th percentiles of distribution times achieved by each system. The results demonstrate that PeerSync significantly outperforms the other systems in real-world conditions.
  • At the 90th percentile (P90), PeerSync achieves an image distribution time of 190.79 seconds, representing a 39.7% improvement over Kraken, a 60.9% improvement over Dragonfly, and a 75.9% improvement over the Baseline.
  • At the 99th percentile (P99), PeerSync maintains its advantage, with a distribution time of 243.89 seconds, which is 27.8% faster than Kraken, 53.9% faster than Dragonfly, and 69.9% faster than the Baseline.
TABLE XI: The 90th and 99th percentile of image distribution times.
Solution Baseline Dragonfly Kraken PeerSync
P90 (s) 790.13 487.50 316.42 190.79
P99 (s) 811.23 528.49 337.75 243.89
The results show that PeerSync performs better than the other methods significantly in our real-world experiments.
V Related Works
Numerous solutions have been explored to enhance distribution speeds in container image delivery. In this section, we briefly review key developments in this area, covering both container-specific solutions and methodologies applied to other resource types.
Before the era of edge computing, several pioneering works focused on fast resource delivery in data center networks. VDN [36] leverages hierarchical network topology and shared chunks between virtual machine images to achieve a 30–80$\times$ speedup for large virtual machine images under heavy traffic in data center networks. Similarly, VMTorrent [37] improves P2P methods used in live streaming, employing block prioritization, profile-based execution prefetch, on-demand fetch, and decoupling of virtual machine image presentation from the underlying data stream. Experiments show VMTorrent can achieve a 30x speedup over traditional network storage.
Since the rise of Docker, containers have become central to deploying workloads in data centers, and distributing container images has emerged as a key issue. While similar to virtual machine images, containers have unique properties, such as a layered structure and registry-based hosting. DevOps practices demand more automated and integrated workload deployment solutions. FID [22], an early P2P-based approach for fast container image distribution, adapts BitTorrent to provide an integrated distribution system. In industry, Dragonfly [2] and Kraken [43] are prominent P2P container image distribution systems. These methods outperform traditional approaches by utilizing spare bandwidth between clients.
However, these solutions do not explicitly address edge computing environments. While Dragonfly [2] and Kraken [43] are P2P-based, they rely on multiple centralized components, which can degrade service quality in the event of component failure. Edge environments are less stable than traditional cloud facilities, and deploying these centralized components in the cloud requires manual configuration across different edge locations. Additionally, edge devices typically have weaker network capabilities, so unrestricted use of P2P protocols can lead to uplink congestion. EdgePier [8], based on IPFS [9], provides a fully decentralized container image distribution system tailored to edge computing. Similarly, Gazzetti et al. [17] proposed a decentralized solution with managers that compute optimal network topologies.
Beyond P2P, another trend involves distributed storage to form a shared storage layer among participating nodes. CoMICon [31], Cider [14], and Wharf [49] use distributed filesystems to accelerate container provisioning. However, these methods are primarily focused on data centers, and implementing distributed storage in edge environments remains challenging [30].
More recently, Starlight [11] has introduced a novel approach by redesigning the container image architecture. It transmits only the necessary components, significantly reducing transfer times. However, Starlight’s reliance on specific container engines limits its applicability, and it continues to use a traditional client-server registry architecture, which restricts throughput under high bandwidth loads. Future work could integrate Starlight with PeerSync to achieve (a) distributed transmission of large data chunks and (b) selective transmission of necessary data, further improving the containerization ecosystem.
VI Concluding Remarks
In this paper, we introduced PeerSync, an innovative fully decentralized image distribution system tailored for containerized model inference workloads at the network edge. PeerSync leverages advanced P2P downloading techniques, outperforming traditional methods by dynamically adapting to fluctuating network conditions and content popularity. Its embedded autonomous tracker enhances resilience by eliminating single points of failure, making it particularly effective in edge environments characterized by unstable networks and resource-constrained devices. Additionally, the cache cleaner efficiently optimizes storage usage while maintaining performance, addressing a key limitation of edge devices.
Future work offers several opportunities to extend PeerSync’s capabilities. a) Dynamic block sizing: While the current static block size approach performs well, incorporating adaptive block sizes with multiple Merkle tree versions could further enhance efficiency with minimal overhead. b) Support for heterogeneous LANs: Adapting to diverse network profiles, such as WLANs, using a combination of user-defined tags and auto-detection techniques could improve performance in non-uniform environments. c) Plugin-based algorithm design: Providing users with greater flexibility through plugin-based customization of algorithms could further expand PeerSync’s adaptability.
Integrating PeerSync with Starlight [11] represents another exciting avenue for exploration. The combination of PeerSync’s decentralized, adaptive framework and Starlight’s innovative protocols, filesystem, and image storage formats could pave the way for next-generation containerization ecosystems, optimized for the evolving model inference demands at the network edge.